# Iterating $\mathsf{PP}$-packages without Choice: A Cohen symmetric seed and a localization framework

### v3

@@ 5.2 Iteration-level diagonal automorphisms
Let $\beta<\lambda$ and let $D\subseteq\mathrm{Coords}_{\beta}$. If $H\in\tilde{\mathscr{F}}^{*}_{\beta}$ contains $\Delta_{\beta}^{\uparrow}(E,D)$, then $(\pi^{\beta}_{\lambda})^{-1}[H]$ contains $\Delta_{\lambda}^{\uparrow}(E,D)$.
@@
Let $g\in\Delta_{\lambda}^{\uparrow}(E,D)$. By Lemma 5.41, $\pi^{\beta}_{\lambda}(g)\in\Delta_{\beta}^{\uparrow}(E,D)\subseteq H$. Hence $g\in(\pi^{\beta}_{\lambda})^{-1}[H]$. ∎
@@
For each $\beta\leq\lambda$, define $D_{\beta}:=D\cap\mathrm{Coords}_{\beta}$, the set of protected coordinates at stages $\leq\beta$.
Key observation from admissibility: Since $(E,D)$ is globally admissible and $\pi\in\operatorname{Fix}(E)$, we have $\operatorname{Fix}(E)\leq\operatorname{Stab}(d)$ for all $d\in D$, hence $\pi\cdot d=d$ for all $d\in D$. In particular, for each successor $\beta+1\leq\lambda$, the set $D_{\beta+1}\setminus D_{\beta}\subseteq D$ is fixed pointwise by $\pi$. This ensures the successor diagonal kernel element $\kappa_{\beta+1}(\pi,D_{\beta+1}\setminus D_{\beta})$ is well-defined (Definition 5.23).
@@
We construct by transfinite recursion a coherent sequence $\langle g_{\beta}:\beta\leq\lambda\rangle$ with $g_{\beta}\in\mathscr{G}_{\beta}$ such that $g_{\beta}=\widehat{\pi}^{\,D_{\beta}}$ as automorphisms of $\mathbb{P}_{\beta}$. More precisely, we maintain the inductive invariant that for all $\gamma<\beta\leq\lambda$, $\pi^{\gamma}_{\beta}(g_{\beta})=g_{\gamma}$. In particular, at a limit stage $\eta$ the family $\langle g_{\beta}:\beta<\eta\rangle$ is coherent, so a direct-limit element exists.
Stage 0: Define $g_{0}:=\pi\in\operatorname{Sym}(\omega_{1})=\mathscr{G}_{0}$. Since $\mathrm{Coords}_{0}=\varnothing$, we have $D_{0}=\varnothing$, so $\widehat{\pi}^{\,D_{0}}=\widehat{\pi}^{\,\varnothing}$, which is the standard $\pi$-action on the Cohen forcing (Definition 5.27, stage-0 clause). Hence $g_{0}=\widehat{\pi}^{\,D_{0}}$.
Successor $\beta+1$: Assume $g_{\beta}\in\mathscr{G}_{\beta}$ with $g_{\beta}=\widehat{\pi}^{\,D_{\beta}}$. Define
$g_{\beta+1}:=\mathrm{spl}(g_{\beta})\cdot\kappa_{\beta+1}\bigl(\pi,\,D_{\beta+1}\setminus D_{\beta}\bigr),$
where $\mathrm{spl}$ is the splitting section from Lemma 5.21 and $\kappa_{\beta+1}$ is the successor diagonal kernel element from Definition 5.23 (applicable by the key observation above).
@@
By Lemma 5.21, $\mathrm{spl}(g_{\beta})\in\mathscr{G}_{\beta+1}$. By Lemma 5.25(i), $\kappa_{\beta+1}(\pi,D_{\beta+1}\setminus D_{\beta})\in\mathscr{G}_{\beta+1}$. Hence $g_{\beta+1}\in\mathscr{G}_{\beta+1}$.
We verify $g_{\beta+1}=\widehat{\pi}^{\,D_{\beta+1}}$ by comparing actions (cf. Definition 5.27):
  • On the $\mathbb{P}_{\beta}$-part: $\mathrm{spl}(g_{\beta})$ acts as $g_{\beta}=\widehat{\pi}^{\,D_{\beta}}$, and $\kappa_{\beta+1}$ acts as identity. The combined action is $\widehat{\pi}^{\,D_{\beta}}$, which agrees with $\widehat{\pi}^{\,D_{\beta+1}}$ restricted to stages $\leq\beta$ (since $D_{\beta+1}\cap\mathrm{Coords}_{\beta}=D_{\beta}$).
  • On new coordinates at stage $\beta+1$: $\mathrm{spl}(g_{\beta})$ acts trivially. The kernel element $\kappa_{\beta+1}(\pi,D_{\beta+1}\setminus D_{\beta})$:
    • acts as $\pi$ on $d\notin D_{\beta+1}\setminus D_{\beta}$, i.e., on new coordinates not in $D$;
    • acts as identity on $d\in D_{\beta+1}\setminus D_{\beta}$, i.e., on new coordinates in $D$.
    This matches Definition 5.27: “act as $\pi$ if $d\notin D$, identity if $d\in D$.”
Limit $\eta$: Define $g_{\eta}$ to be the direct-limit element of $\mathscr{G}_{\eta}$ determined by the family $\langle g_{\beta}:\beta<\eta\rangle$, which is coherent by the inductive invariant above. By the limit-stage group definition ([5, Def. 3.1]), $g_{\eta}\in\mathscr{G}_{\eta}$.
By the uniqueness property of the inverse/direct limit presentation: elements of $\mathscr{G}_{\eta}$ are determined by their projections to all earlier stages ([5, Def. 3.1]). Since $g_{\eta}$ and $\widehat{\pi}^{\,D_{\eta}}$ have identical restrictions to each $\mathbb{P}_{\beta}$ for $\beta<\eta$ (namely, $g_{\beta}=\widehat{\pi}^{\,D_{\beta}}$), we conclude $g_{\eta}=\widehat{\pi}^{\,D_{\eta}}$.
Coherence: We verify $\pi^{\gamma}_{\beta}(g_{\beta})=g_{\gamma}$ for $\gamma<\beta\leq\lambda$.
For successor $\beta+1$:
$\pi^{\beta}_{\beta+1}(g_{\beta+1})$ $=\pi^{\beta}_{\beta+1}\bigl(\mathrm{spl}(g_{\beta})\cdot\kappa_{\beta+1}(\pi,D_{\beta+1}\setminus D_{\beta})\bigr)$
$=\pi^{\beta}_{\beta+1}(\mathrm{spl}(g_{\beta}))\cdot\pi^{\beta}_{\beta+1}(\kappa_{\beta+1}(\ldots))$
$=g_{\beta}\cdot\mathrm{id}=g_{\beta},$
@@
For limits $\eta$, coherence is automatic: $g_{\eta}$ is defined as the direct-limit element, so $\pi^{\gamma}_{\eta}(g_{\eta})=g_{\gamma}$ for all $\gamma<\eta$ by construction.
Conclusion: Taking $\beta=\lambda$, we have $g_{\lambda}=\widehat{\pi}^{\,D_{\lambda}}$. Since $D_{\lambda}=D\cap\mathrm{Coords}_{\lambda}=D$ (since $D\subseteq\mathrm{Coords}_{<\lambda}$ by Definition 5.26(ii)), we have $g_{\lambda}=\widehat{\pi}^{\,D}$ and $g_{\lambda}\in\mathscr{G}_{\lambda}$. Hence $\widehat{\pi}^{\,D}\in\mathscr{G}_{\lambda}$. ∎
@@
Let $\pi\in\operatorname{Fix}(E)$ and let $D^{\prime}\supseteq D$ be such that the diagonal lift $\widehat{\pi}^{\,D^{\prime}}\in\mathscr{G}_{\lambda}$ exists. Then $\widehat{\pi}^{\,D^{\prime}}\in\Delta_{\lambda}^{\uparrow}(E,D)$.
@@
By the characterization definition of $\Delta_{\lambda}^{\uparrow}(E,D)$ (Definition 5.27), we verify:
  1. (i)
    $\pi^{0}_{\lambda}(\widehat{\pi}^{\,D^{\prime}})\in\operatorname{Fix}(E)$: By construction of diagonal lifts (Lemma 5.31(i)), the stage-0 projection of $\widehat{\pi}^{\,D^{\prime}}$ is $\pi$, and $\pi\in\operatorname{Fix}(E)$ by assumption.
  2. (ii)
    $\widehat{\pi}^{\,D^{\prime}}$ fixes every coordinate in $D$ pointwise: By construction (Lemma 5.31(ii)), $\widehat{\pi}^{\,D^{\prime}}$ fixes every coordinate in $D^{\prime}$ pointwise, and $D\subseteq D^{\prime}$.
Hence $\widehat{\pi}^{\,D^{\prime}}\in\Delta_{\lambda}^{\uparrow}(E,D)$. ∎
@@
Let $f$ have countable support $E$ (so $\operatorname{Fix}(E)\leq\operatorname{Stab}(f)$) and let $\dot{s}_{f}$ be the canonical $\mathbb{S}_{[f]}$-name for the $f$-coordinate generic section. Then $\Delta^{\uparrow}(E,\{f\})\leq\operatorname{Stab}(\dot{s}_{f})$, hence $\dot{s}_{f}\in\operatorname{HS}$ (with respect to $\mathscr{F}_{[f]}$).
@@
Let $E\subseteq\omega_{1}$ witness that $f$ has countable support, i.e. $\operatorname{Fix}(E)\leq\operatorname{Stab}(f)$.
Fix any admissible pair $(E^{\prime},D^{\prime})$ such that $E\subseteq E^{\prime}$ and $\{f\}\subseteq D^{\prime}$. Let $\pi\in\operatorname{Fix}(E^{\prime})$. Then the diagonal lift $\hat{\pi}^{D^{\prime}}\in\Delta(E^{\prime},D^{\prime})$ fixes the $f$-coordinate pointwise: for every condition $\vec{p}\in\mathbb{S}_{[f]}$,
$(\hat{\pi}^{D^{\prime}}\cdot\vec{p})(f)=\vec{p}(f).$
Since $\dot{s}_{f}$ is definable (as a $\mathbb{S}_{[f]}$-name) purely from the $f$-coordinate generic section, it follows that $\hat{\pi}^{D^{\prime}}\cdot\dot{s}_{f}=\dot{s}_{f}$. Hence $\Delta(E^{\prime},D^{\prime})\leq\operatorname{Stab}(\dot{s}_{f})$.
As $\Delta^{\uparrow}(E,\{f\})$ is generated by all such subgroups $\Delta(E^{\prime},D^{\prime})$, we have $\Delta^{\uparrow}(E,\{f\})\leq\operatorname{Stab}(\dot{s}_{f})$. Therefore $\operatorname{Stab}(\dot{s}_{f})\in\mathscr{F}_{[f]}$, so $\dot{s}_{f}\in\operatorname{HS}$. ∎
@@
Let $g$ be a package with countable support witnessed by $E\subseteq\omega_{1}$, i.e. $\operatorname{Fix}(E)\leq\operatorname{Stab}(g)$. Let $\dot{r}_{g}$ be the canonical $\mathbb{R}_{[g]}$-name for the $g$-coordinate generic right inverse. Then $\Delta^{\uparrow}(E,\{g\})\leq\operatorname{Stab}(\dot{r}_{g})$, and hence $\dot{r}_{g}\in\operatorname{HS}$ (with respect to $(\mathbb{R}_{[g]},\mathscr{G}_{[g]},\mathscr{F}_{[g]})$).
@@
The proof is identical to Lemma 5.45. Fix an admissible pair $(E^{\prime},D^{\prime})$ with $E\subseteq E^{\prime}$ and $\{g\}\subseteq D^{\prime}$. For $\pi\in\operatorname{Fix}(E^{\prime})$, the diagonal lift $\hat{\pi}^{D^{\prime}}\in\Delta(E^{\prime},D^{\prime})$ fixes the $g$-coordinate pointwise, and $\dot{r}_{g}$ is definable purely from the $g$-coordinate generic object. Hence $\hat{\pi}^{D^{\prime}}\cdot\dot{r}_{g}=\dot{r}_{g}$, so $\Delta(E^{\prime},D^{\prime})\leq\operatorname{Stab}(\dot{r}_{g})$. Therefore $\Delta^{\uparrow}(E,\{g\})\leq\operatorname{Stab}(\dot{r}_{g})$, as required. ∎
Lemma 5.47 (Successor-step HS lift for package generics).
Fix a stage $\beta$ and a package coordinate $d\in\mathrm{NewCoords}_{\beta+1}=\mathrm{Coords}_{\beta+1}\setminus\mathrm{Coords}_{\beta}$ scheduled at stage $\beta$. (e.g. $d=f_{\alpha}$ for a splitting package $\mathbb{Q}_{[f_{\alpha}]}$, or $d=g_{\alpha}$ for a right-inverse package $\mathbb{R}_{[g_{\alpha}]}$). Let $\dot{\tau}$ be a canonical $\dot{\mathbb{Q}}_{\beta}$-name definable purely from the $d$-coordinate generic object (e.g. $\dot{s}_{f_{\alpha}}$ or $\dot{r}_{g_{\alpha}}$).
Assume $d$ has countable support witnessed by $E\subseteq\omega_{1}$ (so $\operatorname{Fix}(E)\leq\operatorname{Stab}(d)$ in $M_{\beta}$). Then the canonical $\mathbb{P}_{\beta+1}$-name $\dot{\tau}^{*}$ is hereditarily symmetric at stage $\beta+1$; in particular, $\dot{\tau}^{*\,G_{\beta+1}}\in M_{\beta+1}\subseteq\mathcal{M}$.
@@
Since $d$ has countable support witnessed by $E$ in $M_{\beta}$, we have $\operatorname{Fix}(E)\leq\operatorname{Stab}(d)$. Thus $(E,\{d\})$ is globally admissible at stage $\beta+1$ (Definition 5.26), so $\Delta_{\beta+1}^{\uparrow}(E,\{d\})$ is one of the $\Delta^{\uparrow}$-generators of $\tilde{\mathscr{F}}^{*}_{\beta+1}$ (Definition 5.48).
Now let $g\in\Delta_{\beta+1}^{\uparrow}(E,\{d\})$. By the characterization in Definition 5.27, $g$ fixes the coordinate $d$ pointwise (Here “fixes the coordinate $d$ pointwise” is exactly the clause in Definition 5.27: the induced automorphism on the $d$-factor forcing ($\dot{\mathbb{Q}}_{d}$ or $\dot{\mathbb{R}}_{d}$, depending on the package) is literally the identity, i.e. $g\cdot\dot{p}=\dot{p}$ for every canonical name $\dot{p}$ for a condition in the $d$-factor.)
Since $\dot{\tau}^{*}$ is definable purely from the $d$-coordinate generic object, it follows that $g\cdot\dot{\tau}^{*}=\dot{\tau}^{*}$. Hence $\Delta_{\beta+1}^{\uparrow}(E,\{d\})\leq\operatorname{Stab}_{\mathscr{G}_{\beta+1}}(\dot{\tau}^{*})$.
Therefore $\operatorname{Stab}_{\mathscr{G}_{\beta+1}}(\dot{\tau}^{*})\in\tilde{\mathscr{F}}^{*}_{\beta+1}$ by upward closure, so $\dot{\tau}^{*}\in\operatorname{HS}_{\beta+1}$ and $\dot{\tau}^{*\,G_{\beta+1}}\in M_{\beta+1}\subseteq\mathcal{M}$. ∎

### v4

@@ 5.2 Iteration-level diagonal automorphisms
Let $\beta<\lambda$, let $E\subseteq\omega_{1}$, and let $D\subseteq\mathrm{Coords}_{\beta}$. If $H\in\tilde{\mathscr{F}}^{*}_{\beta}$ contains $\Delta_{\beta}^{\uparrow}(E,D)$, then $(\pi^{\beta}_{\lambda})^{-1}[H]$ contains $\Delta_{\lambda}^{\uparrow}(E,D)$.
@@
Let $k\in\Delta_{\lambda}^{\uparrow}(E,D)$. By Lemma 5.43, $\pi^{\beta}_{\lambda}(k)\in\Delta_{\beta}^{\uparrow}(E,D)\subseteq H$. Hence $k\in(\pi^{\beta}_{\lambda})^{-1}[H]$. ∎
@@
For each $\beta\leq\lambda$ let $D_{\beta}:=D\cap\mathrm{Coords}_{\beta}$. By Definition 5.28, $\operatorname{Fix}(E)\leq\operatorname{Stab}(d)$ for every $d\in D$, hence $\pi\cdot d=d$ for all $d\in D$. In particular, for each successor $\beta+1\leq\lambda$ the set
$D^{\mathrm{new}}_{\beta+1}:=D_{\beta+1}\setminus D_{\beta}$
is pointwise fixed by $\pi$, so $\kappa_{\beta+1}(\pi,D^{\mathrm{new}}_{\beta+1})$ is well-defined (Definition 5.25).
@@
We build by transfinite recursion a coherent sequence $\langle k_{\beta}:\beta\leq\lambda\rangle$ with $k_{\beta}\in\mathscr{G}_{\beta}$ such that $k_{\beta}=\widehat{\pi}^{\,D_{\beta}}$ as automorphisms of $\mathbb{P}_{\beta}$, and $\pi^{\gamma}_{\beta}(k_{\beta})=k_{\gamma}$ for all $\gamma<\beta\leq\lambda$.
Stage 0. Set $k_{0}:=\pi\in\mathscr{G}_{0}=\mathscr{G}$. Since $D_{0}=\varnothing$, this agrees with $\widehat{\pi}^{\,D_{0}}=\widehat{\pi}^{\,\emptyset}$ on $\mathbb{P}_{0}$.
Successor $\beta+1$. Assume $k_{\beta}\in\mathscr{G}_{\beta}$ with $k_{\beta}=\widehat{\pi}^{\,D_{\beta}}$. Define
$k_{\beta+1}:=\mathrm{spl}(k_{\beta})\cdot\kappa_{\beta+1}\!\left(\pi,\,D^{\mathrm{new}}_{\beta+1}\right),$
@@
where $\mathrm{spl}$ is the splitting section from Lemma 5.21. By Lemma 5.21, $\mathrm{spl}(k_{\beta})\in\mathscr{G}_{\beta+1}$, and by Lemma 5.27(i) we have $\kappa_{\beta+1}(\pi,D^{\mathrm{new}}_{\beta+1})\in\mathscr{G}_{\beta+1}$, hence $k_{\beta+1}\in\mathscr{G}_{\beta+1}$.
Moreover, $k_{\beta+1}$ agrees with $\widehat{\pi}^{\,D_{\beta+1}}$ on $\mathbb{P}_{\beta+1}$: on the $\mathbb{P}_{\beta}$-part, $\mathrm{spl}(k_{\beta})$ acts as $k_{\beta}=\widehat{\pi}^{\,D_{\beta}}$ and the kernel element acts trivially; on the new coordinates at stage $\beta+1$, $\mathrm{spl}(k_{\beta})$ acts trivially and $\kappa_{\beta+1}(\pi,D^{\mathrm{new}}_{\beta+1})$ acts as $\pi$ outside $D^{\mathrm{new}}_{\beta+1}$ and as the identity on $D^{\mathrm{new}}_{\beta+1}$, which is exactly the defining clause of $\widehat{\pi}^{\,D_{\beta+1}}$ (Definition 5.29).
Finally,
$\pi^{\beta}_{\beta+1}(k_{\beta+1})=\pi^{\beta}_{\beta+1}(\mathrm{spl}(k_{\beta}))\cdot\pi^{\beta}_{\beta+1}(\kappa_{\beta+1}(\ldots))=k_{\beta}\cdot\mathrm{id}=k_{\beta},$
@@
Limit $\eta\leq\lambda$. Let $\eta$ be limit and assume $\langle k_{\beta}:\beta<\eta\rangle$ has been constructed coherently. By the limit-stage group definition ([5, Def. 3.1]), there is a unique direct-limit element $k_{\eta}\in\mathscr{G}_{\eta}$ with $\pi^{\beta}_{\eta}(k_{\eta})=k_{\beta}$ for all $\beta<\eta$. Since $\widehat{\pi}^{\,D_{\eta}}$ has the same restrictions to each $\mathbb{P}_{\beta}$ for $\beta<\eta$, we have $k_{\eta}=\widehat{\pi}^{\,D_{\eta}}$.
Conclusion. At $\beta=\lambda$ we obtain $k_{\lambda}=\widehat{\pi}^{\,D_{\lambda}}\in\mathscr{G}_{\lambda}$. Because $D\subseteq\mathrm{Coords}_{<\lambda}$ (Definition 5.28(ii)), we have $D_{\lambda}=D$, hence $\widehat{\pi}^{\,D}\in\mathscr{G}_{\lambda}$. ∎
@@
Let $\lambda$ be a stage, let $E\subseteq\omega_{1}$, and let $D\subseteq\mathrm{Coords}_{<\lambda}$. Let $\pi\in\operatorname{Fix}(E)$ and let $D^{\prime}\supseteq D$ be such that the diagonal lift $\widehat{\pi}^{\,D^{\prime}}\in\mathscr{G}_{\lambda}$ exists. Then
$\widehat{\pi}^{\,D^{\prime}}\in\Delta_{\lambda}^{\uparrow}(E,D).$
@@
Since $\widehat{\pi}^{\,D^{\prime}}\in\mathscr{G}_{\lambda}$ and $\widehat{\pi}^{\,D^{\prime}}$ restricts to $\pi$ on the Cohen base (Lemma 5.33(i)), we have $\pi^{0}_{\lambda}(\widehat{\pi}^{\,D^{\prime}})=\pi\in\operatorname{Fix}(E)$ (e.g. by Corollary 5.34).
Moreover, by Lemma 5.33(ii) the automorphism $\widehat{\pi}^{\,D^{\prime}}$ fixes every coordinate in $D^{\prime}$ pointwise, hence fixes every coordinate in $D\subseteq D^{\prime}$ pointwise. Therefore $\widehat{\pi}^{\,D^{\prime}}\in\Delta_{\lambda}^{\uparrow}(E,D)$ by Definition 5.29(A). ∎
@@
Let $f$ have countable support witnessed by $E\in[\omega_{1}]^{\leq\omega}$ (so $\operatorname{Fix}(E)\leq\operatorname{Stab}(f)$), and let $\dot{s}_{f}$ be the canonical $\mathbb{S}_{[f]}$-name for the $f$-coordinate generic section. Then
$\Delta^{\uparrow}(E,\{f\})\ \leq\ \operatorname{Stab}_{\mathscr{G}_{[f]}}(\dot{s}_{f}),$
and hence $\dot{s}_{f}$ is hereditarily symmetric for the orbit package system $(\mathbb{S}_{[f]},\mathscr{G}_{[f]},\mathscr{F}_{[f]})$.
@@
Fix $E\in[\omega_{1}]^{\leq\omega}$ with $\operatorname{Fix}(E)\leq\operatorname{Stab}(f)$.
Let $(E^{\prime},D^{\prime})$ be any admissible pair in the sense of Definition 5.11 such that $E\subseteq E^{\prime}$ and $f\in D^{\prime}$. Let $\pi\in\operatorname{Fix}(E^{\prime})$ and consider the diagonal-cancellation automorphism $\widehat{\pi}^{\,D^{\prime}}\in\Delta(E^{\prime},D^{\prime})$ (Definition 5.11). Because $f\in D^{\prime}$, this automorphism fixes the $f$-coordinate pointwise, i.e. for every $\vec{p}\in\mathbb{S}_{[f]}$,
$(\widehat{\pi}^{\,D^{\prime}}\cdot\vec{p})(f)=\vec{p}(f).$
Since $\dot{s}_{f}$ is the canonical name definable purely from the $f$-coordinate generic section, we have $\widehat{\pi}^{\,D^{\prime}}\cdot\dot{s}_{f}=\dot{s}_{f}$. Thus $\Delta(E^{\prime},D^{\prime})\leq\operatorname{Stab}_{\mathscr{G}_{[f]}}(\dot{s}_{f})$.
By Definition 5.13, $\Delta^{\uparrow}(E,\{f\})$ is generated by all such $\Delta(E^{\prime},D^{\prime})$ with $E\subseteq E^{\prime}$ and $f\in D^{\prime}$, so $\Delta^{\uparrow}(E,\{f\})\leq\operatorname{Stab}_{\mathscr{G}_{[f]}}(\dot{s}_{f})$. Therefore $\operatorname{Stab}_{\mathscr{G}_{[f]}}(\dot{s}_{f})\in\mathscr{F}_{[f]}$ by Definition 5.13, and $\dot{s}_{f}$ is hereditarily symmetric. ∎
@@
Let $g$ be a package with countable support witnessed by $E\in[\omega_{1}]^{\leq\omega}$, i.e. $\operatorname{Fix}(E)\leq\operatorname{Stab}(g)$. Let $\dot{r}_{g}$ be the canonical $\mathbb{R}_{[g]}$-name for the $g$-coordinate generic right inverse. Then
$\Delta^{\uparrow}(E,\{g\})\ \leq\ \operatorname{Stab}_{\mathscr{G}_{[g]}}(\dot{r}_{g}),$
and hence $\dot{r}_{g}$ is hereditarily symmetric for $(\mathbb{R}_{[g]},\mathscr{G}_{[g]},\mathscr{F}_{[g]})$.
@@
Fix $E\in[\omega_{1}]^{\leq\omega}$ with $\operatorname{Fix}(E)\leq\operatorname{Stab}(g)$.
Let $(E^{\prime},D^{\prime})$ be any admissible pair (Definition 5.11) such that $E\subseteq E^{\prime}$ and $g\in D^{\prime}$. For $\pi\in\operatorname{Fix}(E^{\prime})$, the diagonal-cancellation map $\widehat{\pi}^{\,D^{\prime}}\in\Delta(E^{\prime},D^{\prime})$ fixes the $g$-coordinate pointwise (since $g\in D^{\prime}$). Because $\dot{r}_{g}$ is the canonical name definable purely from the $g$-coordinate generic object, we have $\widehat{\pi}^{\,D^{\prime}}\cdot\dot{r}_{g}=\dot{r}_{g}$. Hence $\Delta(E^{\prime},D^{\prime})\leq\operatorname{Stab}_{\mathscr{G}_{[g]}}(\dot{r}_{g})$.
By Definition 5.13, $\Delta^{\uparrow}(E,\{g\})$ is generated by all such $\Delta(E^{\prime},D^{\prime})$ with $E\subseteq E^{\prime}$ and $g\in D^{\prime}$, so $\Delta^{\uparrow}(E,\{g\})\leq\operatorname{Stab}_{\mathscr{G}_{[g]}}(\dot{r}_{g})$. Therefore $\operatorname{Stab}_{\mathscr{G}_{[g]}}(\dot{r}_{g})\in\mathscr{F}_{[g]}$ (Definition 5.13), and $\dot{r}_{g}$ is hereditarily symmetric. ∎
Lemma 5.49 (Successor-stage diagonal cancellation yields $\operatorname{HS}$-lifts).
Let $\beta$ be an ordinal and consider the successor stage $\beta+1$ with $\mathbb{P}_{\beta+1}\cong\mathbb{P}_{\beta}*\dot{\mathbb{Q}}_{\beta}$. Let $d$ be a package coordinate of $\dot{\mathbb{Q}}_{\beta}$ (e.g. a $\mathbb{Q}$-package coordinate or a $\mathbb{R}$-package coordinate), and let $\dot{\tau}$ be a canonical $\dot{\mathbb{Q}}_{\beta}$-name definable purely from the $d$-coordinate generic object (e.g. $\dot{s}_{f_{\beta}}$ or $\dot{r}_{g_{\beta}}$). Let $\dot{\tau}^{*}$ denote the corresponding canonical $\mathbb{P}_{\beta+1}$-name under this two-step presentation.
Assume $d$ has countable support witnessed by $E\in[\omega_{1}]^{\leq\omega}$ in $M_{\beta}$, so $\operatorname{Fix}(E)\leq\operatorname{Stab}(d)$ in $M_{\beta}$. Then $\dot{\tau}^{*}\in\operatorname{HS}_{\beta+1}$; in particular, $\dot{\tau}^{*\,G_{\beta+1}}\in M_{\beta+1}\subseteq\mathcal{M}$.
@@
From $\operatorname{Fix}(E)\leq\operatorname{Stab}(d)$ we get that $(E,\{d\})$ is globally admissible at stage $\beta+1$ (Definition 5.28). Hence $\Delta_{\beta+1}^{\uparrow}(E,\{d\})$ is one of the designated $\Delta^{\uparrow}$-generators of $\tilde{\mathscr{F}}^{*}_{\beta+1}$ (Definition 5.50).
Let $k\in\Delta_{\beta+1}^{\uparrow}(E,\{d\})$. By Definition 5.29, $k$ fixes the coordinate $d$ pointwise (i.e. its induced action on the $d$-factor forcing is the identity). Since $\dot{\tau}^{*}$ is a canonical name definable purely from the $d$-coordinate generic object, it follows that $k\cdot\dot{\tau}^{*}=\dot{\tau}^{*}$. Therefore
$\Delta_{\beta+1}^{\uparrow}(E,\{d\})\ \leq\ \operatorname{Stab}_{\mathscr{G}_{\beta+1}}(\dot{\tau}^{*}).$
By upward closure of $\tilde{\mathscr{F}}^{*}_{\beta+1}$, we get $\operatorname{Stab}_{\mathscr{G}_{\beta+1}}(\dot{\tau}^{*})\in\tilde{\mathscr{F}}^{*}_{\beta+1}$, hence $\dot{\tau}^{*}\in\operatorname{HS}_{\beta+1}$. The interpretation $\dot{\tau}^{*\,G_{\beta+1}}\in M_{\beta+1}\subseteq\mathcal{M}$ follows. ∎